OpenStudy (anonymous):

a store mixes Kenyan coffee worth 10$ per kilogram and Venezuelan coffee worth 13 $ per kilogram. Find how much each should be used to make a 348 kilogram mixture... I'm not even going to tell you the # I cam up with lol....help

OpenStudy (anonymous):

how much does this mixture cost?

OpenStudy (hoblos):

do we have a specific budget

OpenStudy (anonymous):

does not day just states the total mixture..:(

OpenStudy (anonymous):

nope 10$ for kenyan and 13$ for venex

OpenStudy (anonymous):

well there isn't a unique answer to this. you can make a 50/50 mixture...a 70/30 mixture...anything. there is more to this question than you're posting

OpenStudy (anonymous):

oh sorry sorry yes left out a key component...the mixture is to sell for 11$ per kilogram

OpenStudy (anonymous):

ah much better

OpenStudy (anonymous):

sowwwy <3

OpenStudy (anonymous):

thats why i am here apparently lol

OpenStudy (hoblos):

so the total cost is 348*11 = $3828 x+y = 348 10x + 13y = 3828 can you solve this system?

OpenStudy (mertsj):

10k+13(348-k)=11(348)

OpenStudy (anonymous):

you guys are geniusesssss..hugzz

OpenStudy (mertsj):

we know.

OpenStudy (anonymous):

heehee

OpenStudy (mertsj):

lol